# AdaptInfer: Adaptive Token Pruning for Vision–Language Model Inference with Dynamical Text Guidance

### v1

@@ Introduction
Since both attention matrices are already computed during the forward pass, this additional cost is minimal relative to the main transformer computations.
During the decode stage, the FLOPs of each layer can be estimated by
@@
Layer-wise Pruning Schedule
Following the attention shift analysis described above, we design a pruning schedule that aligns with the attention dynamics observed in VLMs. Our goal is to perform aggressive pruning while maintaining the inference accuracy. To balance the trade-off, we suggest that pruning should avoid the regions with intense attention shifts. Because the probability of incorrect pruning is quite high, tokens that are not important in the past layers are likely to become important in the next layer. On the contrary, pruning just before or after the dense regions can reduce this risk to some extent.
Based on the consistent attention shift distributions across both MME and TextVQA datasets, we select layers 1, 10, and 20 as pruning locations in our framework. Pruning at layer 1 implies performing a forward pass through the first layer, and subsequently using its attention outputs to identify and remove redundant vision tokens based on their ranked importance. Since layer 1 exhibits the earliest dense attention shift, retaining the salient vision tokens at this stage allows them to effectively participate in subsequent information exchange. Layer 10 begins to mark a region of high attention volatility until layer 20. As a result, we preserve the vision tokens within this range, only compress the token numbers at the beginning and the end. At layer 20 and after, where shift activity is negligible, we suggest that the cross-modality information exchange is already sufficient, so we apply aggressive pruning at this layer by removing nearly all remaining visual tokens.
This schedule balances caution and efficiency. Compared to heuristic or uniform pruning schemes, our approach is both data-driven and architecture-aware, requiring no expensive hyperparameter tuning while generalizing well across tasks and datasets.
Discussion
@@
The chosen pruning hyperparameters come from empirical observations on LLaVA-1.5-7B, and thus are tailored specifically to VLMs built upon the LLaMA-7B backbone. Nevertheless, our proposed adaptive pruning schedule is generalizable for it can be easily transferred to other models with different parameter scales or architectures by performing a simple, offline attention shift analysis. The extra experiments will be presented in the technical appendix.
@@
Moreover, one of the datasets used in our observational studies is MME (Fu et al. 2023), a comprehensive multimodal benchmark comprising two major categories and fourteen subcategories. The consistent statistical patterns indicate that the attention dynamics are largely stable and transferable across different types of multimodal tasks.
Experiment
Method Avg. Tokens MME GQA MMBench SQA TextVQA Acc. Ratio (%)
@@
AdaptInfer (Ours) 128 1796 58.5 64.0 70.0 57.0 97.7 ($\downarrow$ 2.3)
@@
AdaptInfer (Ours) 64 1652 53.5 61.3 70.0 54.7 92.9 ($\downarrow$ 7.1)
Table 1: Comparison of methods under different pruning budgets on LLava-1.5-7B. We report results on five datasets, average retained tokens, accuracy, overall ratios and FLOPs. Results in bold present the highest accuracy under same pruning budgets.
Experimental Settings
Datasets
For multimodal evaluation, we test our solutions on five widely used benchmarks, including MME (Fu et al. 2023), GQA (Hudson and Manning 2019), MMBench (Liu et al. 2024b), ScienceQA (SQA) (Lu et al. 2022) and TextVQA (Singh et al. 2019).
Baselines
We select four classic and latest VLM acceleration works in the field of vision token sparsification within the plug-and-play paradigm in LLM forward pass as baselines, including FastV (Chen et al. 2025a), ToMe (Bolya et al. 2023), Pyramid Drop (PDrop) (Xing et al. 2024) and SparseVLM (Zhang et al. 2025). Note that we only use the training-free version of PDrop to fit our requirements.
Implement Details
To ensure a fair and comprehensive comparison between baseline methods and AdaptInfer, we report the results in different average retained token budgets (e.g. 128, 64, 48 and 32). AdaptInfer utilizes the pruning parameters described above while other baselines keep their original settings for all core components. Our experiments are performed on two different types of VLMs, including LLava-1.5-7B (Liu et al. 2023b) and InternVL-chat-7B (Chen et al. 2024a). Our experiments are conducted on a server with 8 $\times$ Nvidia RTX 4090 24G with a cpu of Intel Xeon Platinum 8358P.
Main Results
@@
Tokens MME SQA TextVQA Ratio (%)
@@
Method Tokens Metrics MME GQA MMBench TextVQA Average
@@
AdaptInfer (Ours) 64 FLOPs (T) 0.975 0.959 1.317 1.305 1.139 ($\downarrow$ 74.3%)
Latency (ms) 31.0 30.9 34.5 35.5 33.0 ($\downarrow$ 61.3%)
In Table 1, we report the performance of AdaptInfer on LLava-1.5-7B. We provide comparisons with other baselines under average retained vision token numbers of 128 and 64, with metrics of inference accuracy on each dataset. The accuracy has an upper bound of Vanilla LLava with all 576 tokens kept, so an overall accuracy ratio is supported by each baseline. As shown in this Table, AdaptInfer achieves the highest overall accuracy under both 128 and 64 vision token budgets. Under 128 token budgets, our average accuracy ratio reaches 97.7%, which is 0.9% higher than the second-best method, SparseVLM. While this may seem like a small margin, it is in fact approaching the upper bound of Vanilla LLava. In practice, retaining only 64 tokens on average reduces the prefill computational load by 88.9%, enabling significantly more efficient inference. Despite this aggressive pruning, our framework still preserves 92.9% of the original inference accuracy, and outperforms the second-best method, which scores 91.4%, with a clear improvement of 1.5%. The results highlight the robustness of our approach even under tight token constraints.
In Figure 5, we present the performance trends of AdaptInfer with the latest baseline methods SparseVLM and PDrop. AdaptInfer outperforms others in all retained token budgets with a clear margin.
We further evaluate the proposed AdaptInfer on InternVL‑Chat‑7B (Chen et al. 2024a) in Table 2. InternVL adopts a more powerful 6B parameter ViT encoder, which produces finer‑grained vision tokens with richer hidden states. Experiments confirm that our token‑sparsification strategy preserves the original semantic richness of these vision features with high inference accuracy. Under token budgets of 128 and 64, AdaptInfer maintains a high accuracy ration of 97.7% and 91.9% respectively. However, in the case of extreme pruning, where only 32 average vision tokens are retained, the accuracy has a significant drop, reaching only 81.5%. This is because under such extreme pruning conditions, a large number of informative tokens were inevitably removed, which affected the performance.
Latency Test
We replicate the performance of PDrop (Xing et al. 2024) and SparseVLM (Zhang et al. 2025) under 64-token budget, and compare them with that of AdaptInfer. The comparisons are based on four commonly used benchmarks, including MME, GQA, MMBench and TextVQA. According to the table, while theoretical FLOPs estimations are close, our implementation on AdaptInfer reaches a lower average cuda latency of 33.0 ms per sample than 34.5 ms and 36.7 ms by PDrop and SparseVLM. After all, our method rarely introduces additional computational load, such as any extra attention computation or offline feature matching step.
@@
AdaptInfer 32 1190 49.8 74.6
Ablation Study
Firstly, in order to illustrate the necessity of dynamic text guidance, we provide an experiment to compare results with a static text guidance work SparseVLM in Table 4. Also, we conduct this study as an exploration of extreme pruning. In the experiment, AdaptInfer consistently outperforms SparseVLM on MME and TextVQA (TVQA) under extreme low pruning budgets of only 48 and 32 vision tokens retained averagely. Specifically, AdaptInfer can keep a relatively high overall performance of 85.5% and 74.6% respectively. These results prove the effectiveness of our dynamic text guidance design.
Furthermore, we conduct an intuitive study to assess the effectiveness of our pruning hyperparameters, which are inspired by the observations. We compare our observation-driven pruning strategy with three baselines: uniform pruning, single-layer pruning, and random-layer pruning which averages results over five random configurations. The results shown in Table 5 confirm that the chosen hyperparameters are the best suited for AdaptInfer on the LLaMa-7B backbone.
Method Pruning Loc. MME GQA TextVQA
Ours 1,10,20 1796 58.5 57.0
Uniform 0,9,18,27 1792 57.6 56.2
Uniform 2,12,22 1764 57.5 56.7
Single 1 1647 53.7 56.3
Random - 1693 56.4 56.0
@@
Conclusion
This paper proposes a novel plug-and-play solution AdaptInfer for VLM acceleration via dynamic text-guided pruning. We further provide an offline analysis of cross-attention shifts, which motivates a principled pruning schedule. Our VLM acceleration plugin introduces minimal additional computational overhead while maintaining high accuracy. Specially, our AdaptInfer achieves SOTA accuracy under all the token budgets chosen in the experiments. For instance, AdaptInfer reduces CUDA latency by 61.3% and retains an average accuracy of 92.9% on LLaVA-1.5-7B, with only 64 vision tokens preserved per layer on average.
Appendix
Appendix A Change-Point Detection
To identify shifts in cross-attention across transformer layers, we employ change-point detection (Truong, Oudre, and Vayatis 2020) on cumulative cross-attention curves. The intuition is that a significant change in attention behavior corresponds to a noticeable inflection in the cumulative attention distribution over layers.
@@ Appendix A Change-Point Detection
Appendix B Attention Shift Detection on LLava-1.5-13B
(a) Attention shift points distribution on MME.
(b) Attention shift points distribution on TextVQA.
Figure A.6: Layer-wise distribution of attention shifts on LLava-1.5-13B.
Figure A.7: Visualization of AdaptInfer on different examples.
To validate the transferability of our proposed pruning schedule, we conduct attention shift detection on LLaVA-1.5-13B (Liu et al. 2023b), which is built upon a 40-layer LLaMA-2-13B backbone (Touvron et al. 2023). The resulting distributions on two benchmark datasets, MME and TextVQA, are presented in Figure A.6. We observe that the distribution of attention shift points still remains highly consistent on MME and TextVQA. The early-stage concentration of attention shifts consistently occurs at layer 1. Meanwhile, in later layers, dense change points are slightly delayed compared to those in LLaVA-1.5-7B, which typically occurs around layers 11 to 22. This suggests that for LLaVA-1.5-13B and other VLMs built on the LLaMA-2-13B backbone, an appropriate pruning schedule should select representative layers such as 1, 11, and 22.
Moreover, our attention shift detection pipeline is both simple and lightweight. For example, on the MME dataset, extracting 1000 attention maps and performing change-point detection takes only about 6 minutes in total. This confirms the efficiency and transferability of our pruning schedule across different model scales and architectures.
Appendix C Performance of AdaptInfer on LLava-1.5-13B
Using the parameters above, we test our AdaptInfer on LLava-1.5-13Be to further validate the effectiveness of the proposed pruning schedule. The experiments are based on three widely used dataset MME (Fu et al. 2023), MMBench (MMB) (Liu et al. 2024b), and TextVQA (Singh et al. 2019) under 128 and 64 vision token budgets. The results are shown in Table A.6.
@@
128 1803 59.9 68.0 98.6
64 1759 57.9 65.7 95.6
Table A.6: Performance of AdaptInfer on LLava-1.5-13B. Our method still maintain high overall accuracy.
Under the more demanding 13-billion-parameter setting of LLaVA-1.5, AdaptInfer continues to deliver high performance. As Table A.6 shows, pruning the vision token numbers from 576 to 128 on average per layer results in only a 1.4% drop in overall accuracy (from 100% to 98.6%). Even with an aggressive 64 vision token budget, AdaptInfer achieves an 8.9× reduction in vision-token count, the model still retains 95.6% of its full-token accuracy. These numbers demonstrate that our token-adaptive pruning solution scales gracefully with model size, confirming AdaptInfer’s robustness.
Appendix D Visualization

### v2

@@ 3.2.2 Analysis of Computational Complexity
Since both attention matrices are already computed during the forward pass, this additional cost is minimal relative to the main transformer computations. Then, during the decode stage, the FLOPs of each layer can be estimated by
@@
3.2.3 Layer-wise Pruning Schedule
Following the attention shift analysis described above, we design a pruning schedule that aligns with the attention dynamics observed in VLMs. Pruning visual tokens introduces an inherent trade-off between pruning safety and computational savings. To prune aggressively, one must prune early; yet pruning too early inevitably risks removing tokens whose importance has not yet emerged.
Prior works rely directly on attention scores to rank tokens (Xing et al., 2024; Zhang et al., 2025; Chen et al., 2025a). However, we argue that attention scores from all transformer layers are not equally reliable for serving as evidence for pruning. Our attention shift can indicate how reliable the attention-based rankings are across layers. Our attention-shift analysis helps to reveal two schemes: Firstly, in the high-frequency regions (e.g., layer 1 and layers 10–20), token importance is actively being reassigned (either becoming important for the first time or being used up by the model). Attention rankings here are unstable and thus unsafe for pruning. Secondly, in the stable regions (e.g., layers 2–9 and 20+), importance rankings remain consistent and therefore provide reliable pruning signals.
Based on the discussion above, we select to prune vision tokens after layer 1 and after layer 20 on Llava-1.5-7B. Both locations mark the beginning of each stable region, enabling early pruning to save more FLOPs relying on stable attention rankings. To prevent over-pruning at layer 1, which would remove too many informative tokens too early, we choose an additional pruning location between layers 1 and 20. This step is roughly to be placed at layer 10, the midway of layer 1 and 20, which ensure that (1) sufficient depth for each pruning to take effect and (2) little interference with the high-volatility band. This schedule balances caution and efficiency. Compared to heuristic or uniform pruning schemes, our approach is both data-driven and architecture-aware, requiring no expensive hyperparameter tuning while generalizing well across tasks and datasets.
3.3 Discussion
@@
The chosen pruning hyperparameters come from empirical observations on LLaVA-1.5-7B, and thus are tailored specifically to VLMs built upon the LLaMA-7B backbone. Nevertheless, our proposed adaptive pruning schedule is generalizable for it can be easily transferred to other models with different parameter scales or architectures by performing a simple, offline attention shift analysis. Additional experiments on LLaVA-1.5-13B and Qwen2-VL-2B (Wang et al., 2024a) are presented in the Appendix C and D. Based on the attention shift distrubations, we select pruning locations at layer 1, 11 and 22 on LLaVA-1.5-13B, and at layer 0, 9 and 19 on Qwen2-VL-2B for later experiments.
@@
Moreover, one of the datasets used in our observational studies is MME (Fu et al., 2023), a comprehensive multimodal benchmark comprising two major categories and fourteen subcategories. The consistent statistical patterns indicate that the attention dynamics are largely stable and transferable across different types of multimodal tasks.
4 Experiment
4.1 Experimental Settings
4.1.1 Datasets
For multimodal evaluation, we test our solutions on five widely used benchmarks, including MME (Fu et al., 2023), GQA (Hudson & Manning, 2019), MMBench (MMB) (Liu et al., 2024b), ScienceQA (SQA) (Lu et al., 2022), TextVQA (TVQA) (Singh et al., 2019), and POPE (Li et al., 2023b). These datasets together provide a comprehensive evaluation, including vision-question-answering (VQA), optical character recognition (OCR), perception, reasoning, factual grounding and so on.
In addition, we also test our solutions on three popular video-based multimodal benchmarks, TGIF-QA (TGIF) (Jang et al., 2019), MSVD-QA (MSVD) and MSRVTT-QA (MSRVTT) (Xu et al., 2017).
4.1.2 Baselines
We select four classic and latest vision token sparsification frameworks for VLM acceleration within the plug-and-play paradigm in LLM forward pass as baselines, including FastV (Chen et al., 2025a), ToMe (Bolya et al., 2023), Pyramid Drop (PDrop) (Xing et al., 2024) and SparseVLM (Zhang et al., 2025). For PDrop, we only adopt the training-free version of PDrop to fit our requirements. Note that, methods performing token pruning or merging anywhere other than the prefill stage of language models, like vision encoders or the decode stage (Yang et al., 2024; Tong et al., 2025), are not included for comparison, since such approaches are orthogonal to ours and can be used together.
4.1.3 Implement Details
To ensure a fair and comprehensive comparison between baseline methods and AdaptInfer, we report the results in different average retained token budgets (e.g. 128, 64, 48 and 32). AdaptInfer utilizes the pruning parameters described above while other baselines keep their original settings for all core components. Our experiments are performed on two different types of VLMs, including LLava-1.5-7B (Liu et al., 2023b) and InternVL-chat-7B (Chen et al., 2024a). Additional experiments on LLava-1.5-13B are presented in appendix C.2.
4.2 Main Results
Table 1: Comparison of methods under different pruning budgets on LLava-1.5-7B. We report results on five datasets, including average retained tokens, accuracy scores and average ratios. Results in bold present the highest accuracy under same pruning budgets.
Method Avg. Tokens MME GQA MMB SQA TVQA Ratio (%)
@@ 4.2 Main Results
AdaptInfer (Ours) 128 1794 58.5 63.8 69.9 56.8 97.5 ($\downarrow$ 2.5)
@@
AdaptInfer (Ours) 64 1684 53.2 61.7 69.9 54.3 93.1 ($\downarrow$ 6.9)
@@
Tokens MME SQA TVQA Ratio (%)
@@
In Table 1, we report the performance of AdaptInfer on LLava-1.5-7B. We provide comparisons with other baselines under average retained vision token numbers of 128 and 64. The accuracy scores has an upper bound of Vanilla LLava with all 576 tokens kept, so an average accuracy ratio is supported by each baseline. As shown in this Table, AdaptInfer achieves the highest overall accuracy scores under both 128 and 64 vision token budgets. Under 128 token budgets, our average ratio reaches 97.5%, which is 0.7% higher than the second-best method, SparseVLM. While this may seem like a small margin, it is in fact approaching the upper bound of Vanilla LLava. In practice, retaining only 64 tokens on average reduces the prefill token load by 88.9%, enabling significantly more efficient inference. Despite this aggressive pruning, our framework still preserves 93.1% of the original inference accuracy and outperforms SparseVLM, which scores 91.4%, with a clear improvement of 1.7%. In Figure 3, we present the performance trends of AdaptInfer with the latest baseline methods SparseVLM and PDrop. AdaptInfer outperforms others with clear margins.
We further evaluate the proposed AdaptInfer on InternVL‑Chat‑7B (Chen et al., 2024a) in Table 2. InternVL shares the same LLaMA architecture but adopts a more powerful 6B parameter ViT encoder, which produces finer‑grained vision tokens with richer hidden states. Experiments confirm that our token‑sparsification strategy preserves the original semantic richness of these vision features with high inference accuracy. Under token budgets of 128 and 64, AdaptInfer maintains a high accuracy ratio of 97.7% and 91.9% respectively. However, in the case of extreme pruning, where only 32 average vision tokens are retained, the accuracy has a significant drop to only 81.5%. This is because the inevitable removal of a large number of informative tokens affects the performance.
4.3 Evaluation on Qwen2-VL
Qwen2-VL (Wang et al., 2024a) is a more recent and robust VLM series that not only support image-based tasks, but also support video question answering (video QA) tasks. To provide a boarder evaluation, we test AdaptInfer on 8 benchmarks, including 3 video QA datasets (TGIF-QA, MSRVTT-QA, MSVD-QA). Unlike LLaVA, Qwen2-VL dynamically adjusts its vision token counts. Thus, for fairness, we compare SparseVLM and AdaptInfer under the same retained-token ratios (50%, 30%, 10%). The results are summarized in Table 3. Across both image-based QA and multi-frame video QA, AdaptInfer consistently matches or outperforms SparseVLM. Our method performs better particularly under the 30% and 10% retained-token settings, where the pruning becomes more aggressive.
Interestingly, at the 50% retained ratio, both SparseVLM and AdaptInfer slightly outperform the original model. We believe this reflects a widely observed property in multimodal QA: the presence of a large number of redundant visual tokens. Removing these distractors not only accelerates inference but can also improve QA performance by reducing noise. This further confirms the strong practical value of token pruning. More implementation details of AdaptInfer on Qwen2-VL-2B are disclosed in Appendix I.4.
Table 3: Performance of AdaptInfer on Qwen2-VL-2B. Result in bold indicate higher.
Methods Tokens Image-Based Video-Based Ratio
MME TVQA GQA MMB POPE TGIF MSRVTT MSVD
Vanilla 100% 1901 77.8 60.4 71.7 86.9 9.9 30.9 42.3 100%
SparseVLM 50% 1900 76.9 59.9 71.0 86.4 11.7 31.2 42.9 102.1%
AdaptInfer 50% 1909 76.6 60.0 72.1 86.7 11.1 30.5 43.6 101.7%
SparseVLM 30% 1867 73.6 57.5 70.1 84.6 9.3 29.5 41.8 96.4%
AdaptInfer 30% 1885 73.9 58.5 71.4 85.8 10.4 30.2 43.5 99.4%
SparseVLM 10% 1460 51.3 40.4 57.3 40.3 5.0 28.1 30.2 68.6%
AdaptInfer 10% 1721 62.3 48.8 63.7 71.3 5.9 28.2 41.4 81.6%
4.4 Latency Test
We replicate the performance of PDrop (Xing et al., 2024) and SparseVLM (Zhang et al., 2025) under 64-token budget, and compare them with that of AdaptInfer. The comparisons are based on four commonly used benchmarks, including MME, GQA, MMBench (MMB) and TextVQA (TVQA). According to the table, while theoretical FLOPs estimations are close, our implementation on AdaptInfer reaches a lower average cuda latency of 33.0 ms per sample than 34.5 ms and 36.7 ms by PDrop and SparseVLM. After all, our method rarely introduces additional computational load, such as any extra attention computation or offline feature matching step.
Method Tokens Metrics MME GQA MMB TVQA Average
@@ 4.4 Latency Test
AdaptInfer 64 FLOPs (T) 0.975 0.959 1.317 1.305 1.139 ($\downarrow$ 74.3%)
    (Ours) Latency (ms) 31.0 30.9 34.5 35.5 33.0 ($\downarrow$ 61.3%)
4.5 Ablation Study
Firstly, in order to illustrate the necessity of dynamic text guidance, we provide an experiment to compare results with a static text guidance work SparseVLM in Table 5. Also, we conduct this study as an exploration of extreme pruning. In the experiment, AdaptInfer consistently outperforms SparseVLM on MME and TextVQA (TVQA) under extreme low pruning budgets of only 48 and 32 vision tokens retained averagely. Specifically, AdaptInfer can keep a relatively high overall performance of 85.4% and 74.2% respectively. These results prove the effectiveness of our dynamic text guidance design.
Furthermore, we conduct an intuitive study to assess the effectiveness of our pruning hyperparameters. We compare our observation-driven pruning strategy with three baselines: uniform pruning, single-layer pruning, and random-layer pruning which averages results over five random configurations. The results shown in Table 6 confirm that the chosen hyperparameters are the best suited for AdaptInfer on the LLaMa-7B backbone.
@@ 4.5 Ablation Study
AdaptInfer 32 1190 49.3 74.2
@@
Method Pruning Loc. MME MMB TVQA
Ours 1,10,20 1794 63.8 56.8
Uniform 0,9,18,27 1788 62.8 56.4
Uniform 2,12,22 1758 63.2 56.5
Single 1 1668 60.8 56.3
Random - 1679 62.1 55.8
5 Conclusion
This paper proposes a novel plug-and-play solution AdaptInfer for VLM acceleration via dynamic text-guided pruning. We further provide an offline analysis of cross-attention shifts, which motivates a principled pruning schedule. Our VLM acceleration plugin introduces minimal additional computational overhead while maintaining high accuracy. Specially, our AdaptInfer achieves SOTA accuracy under all the token budgets chosen in the experiments. For instance, AdaptInfer reduces CUDA latency by 61.3% and retains an average accuracy of 93.1% on LLaVA-1.5-7B, with only 64 vision tokens preserved per layer on average.
Appendix A Appendix and LLM Usage statement
The chapters below are all technical appendix.
The authors declear that they use LLM tools for grammar refinement and language polishing only. The authors take full responsibility for the entire content of this manuscript.
Appendix B Change-Point Detection
To identify shifts in cross-attention across transformer layers, we employ change-point detection (Truong et al., 2020) on cumulative cross-attention curves. The intuition is that a significant change in attention behavior corresponds to a noticeable inflection in the cumulative attention distribution over layers.
@@ Appendix B Change-Point Detection
Appendix C Experiments on LLaVA-1.5-13B
C.1 Attention Shift Detection on LLaVA-1.5-13B
(a) Attention shift distribution on MME.
(b) Attention shift distribution on TextVQA.
Figure 4: Layer-wise distribution of attention shifts on LLava-1.5-13B.
To validate the transferability of our proposed pruning schedule, we conduct attention shift detection on LLaVA-1.5-13B (Liu et al., 2023b), which is built upon a 40-layer LLaMA-2-13B backbone (Touvron et al., 2023). The resulting distributions on two benchmark datasets, MME and TextVQA, are presented in Figure 4. We observe that the distribution of attention shift points still remains highly consistent on MME and TextVQA. The early-stage concentration of attention shifts consistently occurs at layer 1. Meanwhile, in later layers, dense change points are slightly delayed compared to those in LLaVA-1.5-7B, which typically occurs around layers 11 to 22. This suggests that for LLaVA-1.5-13B and other VLMs built on the LLaMA-2-13B backbone, an appropriate pruning schedule should select representative layers such as 1, 11, and 22.
Moreover, our attention shift detection pipeline is both simple and lightweight. For example, on the MME dataset, extracting 1000 attention maps and performing change-point detection takes only about 8 minutes in total. This confirms the efficiency and transferability of our pruning schedule across different model scales and architectures.
C.2 Performance of AdaptInfer on LLava-1.5-13B
Using the parameters above, we test our AdaptInfer on LLava-1.5-13B to further validate the effectiveness of the proposed pruning schedule. The experiments are based on three widely used dataset MME (Fu et al., 2023), MMBench (MMB) (Liu et al., 2024b), and TextVQA (Singh et al., 2019) under 128 and 64 vision token budgets. The results are shown in Table 7.
@@
128 1813 59.8 67.4 98.5
64 1735 58.1 66.2 95.5
Table 7: Performance of AdaptInfer on LLava-1.5-13B. Our method still maintain high overall accuracy.
Under the more demanding 13-billion-parameter setting of LLaVA-1.5, AdaptInfer continues to deliver high performance. As Table 7 shows, pruning the vision token numbers from 576 to 128 on average per layer results in only a 1.5% drop in overall accuracy (from 100% to 98.5%). Even with an aggressive 64 vision token budget, AdaptInfer achieves an 8.9× reduction in vision-token count, the model still retains 95.5% of its full-token accuracy. These numbers demonstrate that our token-adaptive pruning solution scales gracefully with model size, confirming AdaptInfer’s robustness.
Appendix D Attention Shift Detection on Qwen2-VL-2B
We also conduct our attention shift detection on a different VLM framework, Qwen2-VL-2B (Wang et al., 2024a), which contains 28 transformer layers in its LLM backbone. The attention shift distribution on MME is shown in Fig 5. The high-frequency regions are among layer 10-19, where stable regions are among layer 0-9 and 20-28. Based on our principled pruning schedule, we select the token pruning locations at layer 0, 9, and 19. Moreover, conducting the attention shift analysis on Qwen2-VL- 2B takes only about 5 minutes in total, which illustrates the lightweightness and transferability of the proposed attention shift analysis.
Figure 5: Layer-wise distribution of attention shifts on Qwen2-VL-2B on MME.
Appendix E Pruning Tolerance
To further understand the vision token pruning, we conduct a pruning-tolerance evaluation. Specifically, we apply AdaptInfer to Qwen2-VL-2B and report the performance across 14 multimodal tasks in the MME benchmark, comparing the scores with and without pruning. The results presented in Table 8 indicate that code reasoning, count and color-based tasks have a lower pruning tolerance relatively. Meanwhile, calculation and OCR tasks outperforms the original Qwen due to higher pruning tolerance and the noise reduction effect by token pruning.
Table 8: Pruning-tolerance evaluation on MME.
Task Qwen2-VL-2B AdaptInfer (30% token retained) Ratio
Existence 200 200 100%
Count 145 135 93.1%
Position 158 158 100%
Color 190 180 94.7%
Posters 125 123 98.4%
Celebrity 138 136 98.6%
Scene 159 159 100%
Landmark 174 172 98.9%
Artwork 135 135 100%
OCR 88 103 117%
Commonsense 112 111 99.1%
Calculation 23 33 143%
Translation 163 155 95.1%
Code Reasoning 93 85 91.4%
Appendix F Instance-Level Decision Analysis
There is an inherent trade-off between efficiency and performance in vision token pruning. we admit that aggressive token pruning can induce instance-level decision changes. To examine this effect, we perform a simple instance-level analysis comparing the vanilla Qwen2-VL-2B and 30% token retained AdaptInfer on MME. For each example, we record whether the prediction is (correct/error) before and after pruning. We tabulate the transition, and report the change ratios in Table 9. We find that the vast majority of predictions remain unchanged, while the fraction of Correct to Error cases is small and comparable to the number of Error to Correct cases.
Table 9: Instance-Level change before and after pruning.
Correct to Correct Correct to Error Error to Correct Error to Error
Percentage 78.2% 1.7% 1.3% 18.8%
Appendix G Visualization
Figure 6: Visualization of AdaptInfer on different examples.